Doctors discover more health benefits to vitamin c (found in oranges)? The equilibrium price will (blank) and the equilibrium quantity will (blank) ?
1
Expert's answer
2019-02-14T10:38:07-0500

The recommended daily intake of vitamin C is 90 mg for men and 75 mg for women. Teenagers need from 65 to 75 mg of vitamin C per day, and children need about 35-50 mg. About 3000 mg (or 3 g) of ascorbic acid can be consumed without harm to health - which is equivalent to a kilogram of wild rose berries or 6 kg of oranges or lemons.

Make up the proportion for orange daily allowance (let the average rate of vitamin Bude 80 mg)

3000 mg - 6 kg

80 mg -?

(80 * 6) / 3000 = 0.16 (kg)

If a kilogram of oranges cost $ 1.2, then

The cost of the daily rate of oranges 1.2 * 0.16 = $ 0.192

the cost of the maximum portion of orange will be equal to 1.2 * 6 = $ 72